From GlueXWiki
Jump to: navigation, search

July 7, 2011 FDC meeting


  1. Production Construction Tracking (Dave)
    • Status
    • Second cell tests
    • Other
  2. Engineering (Bill)
    • Package flatness
    • Other: gusset rings, strongback, position measurements
  3. Electronics update (Chris)
  4. Chamber testing at EEL126 (Lubomir)
    • First cell status
  5. Other


Participants: Eugene, Bill, Dave, Chris, Beni, Simon, Caleb, and Lubomir.


- Dave: see the construction tracking link above. For cell#3: gluing the rigid-flexes on the cathodes, some delay here was due problems with the connectors on three of the rigid-flexes and also we run out of tinned flexes and Casey had to tin the un-tinned. We continued with the wire frame#3, we postponed the production for a week since we were not sure about the problem with the frame deformation. The wires were taped and position measurements were done with the new code. Caleb reported that the number of field wires that can't be find increased from the last week. Techs are working on the cell#4 cathodes.

- Lubomir on second cell tests: it was conditioned over the holidays by Anatoly (it took a week in total) and now it holds 2200V, but still tripping every 2-3 hours. The trips might be due to the lexan sheet on the top. Bill proposed to make it conductive for the future tests. The current is very low, 140nA for the whole chamber, especially compared to the first chamber that in similar conditions had ~500nA on one sector and several uA on the rest.

- To find the origin of the current and trips, we checked separately on another laminated wire frame: (1) the current through the HV boards and found to be very small (so the Rohacell is not contributing to the current!) (2) the current through the HV capacitors on the signal side: this current turned out to be significant, between 10 nA and few hundred nA at 2200V. For 96 capacitors one would expect > 1 uA. We will continue investigating the capacitors.


- Survey of the first package showed out of plane deviations up to 5.5mm in horizontal position and that effect goes with the gravity, and ~1mm in vertical position. This explains the top/bottom asymmetry results.

- Bill explained the current design of the connections between the packages. We exchanged different ideas how to use this construction to correct for the out-of-plane deviations of the packages.

- One of the gusset rings is ready and we can start re-assembling the second cell. A survey of the strong-back for the wire frame showed sagging of ~2mm. Bill says it is easy to correct it and it will do it after we are done with the third wire frame. We discussed also what do we need to start the wire deadening.


- Chris: 80 FDC type-1 cables were tested, found one bad. Another 130 FDC cables came, not tested yet. - Eugene: all the ASICS will be ready by the end of the month, but making the pre-amp cards will take many months.

Chamber testing

- As before, we can use one out of four HV sectors. After having understood the effects of the frame deformation, we are now studying the chamber resolution. So far it's not as good as on prototype, especially on the top strips. However, by swapping the cards at the detector it turned out that the effect goes with the card. So, the problem could be in the cards, cables, fADC, or the analysis code.


- Eugene asked Dave how we are the production schedule. Dave: we are late by about a month but expect the be on track for the first package.